OpenStudy (anonymous):

Find the gradient of the function h(x) = square root of (3x^2+2x) at the point where x=2

OpenStudy (anonymous):

\[h(x)=\sqrt{3x^2+2x}\\ h'(x)=\frac{6x+2}{2\sqrt{3x^2+2x}}\\ h'(x)=\frac{3x+1}{\sqrt{3x^2+2x}}\] Find \(h'(2)\).